# Angular momentum of vacuum bubbles in a first-order phase transition

### v1

@@ 6 Conclusions
We worked in the uniform Hubble gauge in tracking the evolution of cosmological perturbations during the FOPT, and numerically obtained the density, velocity, and metric perturbations for each Fourier mode within a chosen range of comoving wavenumbers. Carrying out this task requires knowledge of the evolution of the sound speed and equation of state of the tightly coupled dark fluid. In general, it is possible that the perturbations in the dark sector fluid may deviate from those of a pure radiation fluid, in cases where the equation of state and sound speed squared differ from $1/3$. In particular, we have found that the sound speed squared decreases around the time when the false vacuum fraction has a noticeable drop, takes a value smaller than $1/3$ at percolation, and then approaches the sound speed squared in the true vacuum. For subhorizon modes that have sufficiently short wavelengths, the density perturbations at the percolation time are enhanced relative to the ones in the case of density perturbations evolving in a pure radiation background. In contrast, velocity perturbations are suppressed. In the limit of extremely short-wavelength modes, the density and velocity perturbations at percolation are quenched to values that are determined by the sound speed at percolation. Once the perturbations at the percolation time are known, we calculate the RMS value of the angular momentum and spin. We performed this through phase-space integration, over pairs of Fourier modes, of a product of density and velocity perturbations weighted by a form factor associated with the size of the spherical false vacuum bubble.
@@
From a classical mechanics standpoint, the RMS spin of the false vacuum bubble is dependent on the equatorial velocity, radius, and mass of the sphere. We have chosen the radius of the FV bubble to be the maximal volume at the percolation time, within which no true vacuum bubble can nucleate. We have found a simple scaling relation, which relates the RMS spin with the inverse FOPT time scale $\beta_{*}/H_{*}$, the wall velocity, the degrees of freedom in the dark and SM sectors, and the initial dark-to-SM temperature ratio $r_{T\rm,c}$. In particular, the RMS spin is mainly determined by $(\beta_{*}/H_{*})/v_{\rm w,*}$, $r_{T\rm,c}$, and $T_{*}$. The RMS spin scales as $(\beta_{*}/H_{*})^{2}/v_{\rm w,*}^{2}$, and is inversely proportional to $r_{T\rm,c}^{4}$. The quantity $(\beta_{*}/H_{*})/v_{\rm w,*}$ is mainly determined by the particular FOPT scenario. The critical temperature $T_{\rm c}$ also plays a role in determining the RMS spin, through the direct proportionality between $s_{\rm rms,*}$ and $g_{\rho,{\rm SM}}$. We have found that the RMS spin value can take a wide range of values. For our sample scans with $T_{\rm c}$ between $10\,\mathrm{keV}$ and $100\,\mathrm{GeV}$, and $r_{T\rm,c}$ ranging from 0.1 to 0.4, it can be as low as $\mathcal{O}(10^{-3})$ to a few $\mathcal{O}(10^{3})$. The upper limit on $r_{T\rm,c}$ arises from the Planck constraint on the effective number of relativistic degrees of freedom, $N_{\rm eff}$, and the lower limit on $r_{T\rm,c}$ ensures that an FOPT occurs. A spin larger than unity does not pose any issue, since the false vacuum bubbles are not black holes.
@@ Acknowledgments
We thank Pin-Jung Chen for collaboration during the early stages of this work. JTA thanks Reginald Bernardo, Daniel Harsono, Mehrdad Mirbabayi, Martin Spinrath, Yuhsin Tsai, and Ian Vega for fruitful discussions and insights. JTA and PYT acknowledge the kind support of the National Science and Technology Council of the Republic of China (formerly the Ministry of Science and Technology), with grant number NSTC 111-2811-M-007-018-MY2. DM acknowledges financial support from the U.S. Department of Energy under Grant No. DE-SC0010504. This work used high-performance computing facilities operated by the Center for Informatics and Computation in Astronomy (CICA) at the National Tsing Hua University. This equipment was funded by the Ministry of Education of Taiwan, the National Science and Technology Council of Taiwan, and National Tsing Hua University.

### v2

@@ 6 Conclusions
We worked in the uniform Hubble gauge to track the evolution of cosmological perturbations during the FOPT, and numerically obtained the density, velocity, and metric perturbations for each Fourier mode within a chosen range of comoving wavenumbers. Carrying out this task requires knowledge of the evolution of the sound speed and equation of state of the tightly coupled dark fluid. In general, it is possible that the perturbations in the dark sector fluid may deviate from those of a pure radiation fluid, in cases where the equation of state and sound speed squared differ from $1/3$. In particular, we find that the sound speed squared decreases around the time when the false vacuum fraction has a noticeable drop, takes a value smaller than $1/3$ at percolation, and then approaches the sound speed squared in the true vacuum. In particular, there is a short period in which the sound speed squared becomes negative. Thus, subhorizon modes that have sufficiently short wavelengths will be exponentially enhanced. Because of this enhancement, the mean squared value of the density contrast diverges. This signifies a natural threshold scale, above which the density contrast is sufficiently large to form black holes. Once the perturbations at the percolation time are known, we calculate the RMS value of the angular momentum and spin. We performed this through phase-space integration, over pairs of Fourier modes, of a product of density and velocity perturbations weighted by a form factor associated with the size of the spherical false vacuum bubble. The phase-space integration over Fourier modes is limited to a scale corresponding to the radius of the FV bubbles, and we consider scenarios where this critical scale is at most twice the threshold scale for black hole production.
@@
From a classical mechanics standpoint, the RMS spin of the false vacuum bubble is dependent on the equatorial velocity, radius, and mass of the sphere. We chose the radius of the FV bubble to be the maximal volume at the percolation time, within which no true vacuum bubble can nucleate. We have found a simple scaling relation, which relates the RMS spin with the inverse FOPT time scale $\beta_{*}/H_{*}$, the wall velocity, the degrees of freedom in the dark and SM sectors, and the initial dark-to-SM temperature ratio $r_{T\rm,c}$. In particular, the RMS spin is mainly determined by $(\beta_{*}/H_{*})/v_{\rm w,*}$, $r_{T\rm,c}$, and $T_{*}$. The RMS spin scales as $(\beta_{*}/H_{*})^{2}/v_{\rm w,*}^{2}$ and is inversely proportional to $r_{T\rm,c}^{4}$. The quantity $(\beta_{*}/H_{*})/v_{\rm w,*}$ is mainly determined by the particular FOPT scenario. The critical temperature $T_{\rm c}$ also plays a role in determining the RMS spin through the direct proportionality between $s_{\rm rms,*}$ and $g_{\rho,{\rm SM}}$. We have found that the RMS spin value can take a wide range of values. For our sample scans with $T_{\rm c}$ between $10\,\mathrm{keV}$ and $100\,\mathrm{GeV}$, and $r_{T\rm,c}$ ranging from 0.1 to 0.4, it can be as low as $\mathcal{O}(10^{-5})$ to a few $\mathcal{O}(10)$. The upper limit on $r_{T\rm,c}$ arises from the Planck constraint on the effective number of relativistic degrees of freedom, $N_{\rm eff}$, and the lower limit on $r_{T\rm,c}$ ensures that an FOPT occurs. A spin larger than unity does not pose any issue, since the false vacuum bubbles are not black holes.
@@ Acknowledgments
We thank Pin-Jung Chen for collaboration during the early stages of this work. JTA thanks Reginald Bernardo, Kuan-Yen Chou, Daniel Harsono, Mehrdad Mirbabayi, Michiyasu Nagasawa, Martin Spinrath, Yuhsin Tsai, and Ian Vega for fruitful discussions and insights. JTA and PYT acknowledge the kind support of the National Science and Technology Council of the Republic of China (formerly the Ministry of Science and Technology), with grant number NSTC 111-2811-M-007-018-MY2. DM acknowledges financial support from the U.S. Department of Energy under Grant No. DE-SC0010504. This work used high-performance computing facilities operated by the Center for Informatics and Computation in Astronomy (CICA) at the National Tsing Hua University. This equipment was funded by the Ministry of Education of Taiwan, the National Science and Technology Council of Taiwan, and National Tsing Hua University. We also thank the National Institute of Physics, UP Diliman for allowing us to use the in-house high performance computing cluster.